I'm not exactly sure if this is a Travel and distance problem but that's what I think...
A cyclist rode the 1st 20 mile portion of his workout at a constant speed. He then rode a 16 mile cool down and reduced his speed by 2 miles per hour. Each portion of his workout took the same amount of time. Find the speed during the 1st portion and the speed during the cool down.

Answer by checkley77(12569)

t=x/20
t=(x-2)/16
x/20=(x-2)/16
20(x-2)=16x
20x-40=16x
20x-16x=40
4x=40
x=10 mph is the workout speed
10-2=8 mph is the cool down speed.